Marcus faces a progressive federal income tax structure that has the following marginal tax rates:0 percent on the first $10,000, 10 percent on the next $10,000, 15 percent on the next $10,000 , 25 percent bin the next $10,000 , and 50 percent on all additional income. In the additional , he must pay 5 percent of his income in provincial income tax and 15.3 percent on his Labour income in federal payroll taxes .Marcus earns $60,000 per year in salary and another $10,000 per year in nonlabour income.what is his average tax rate?

a. 17.19 percent 

b. 46.69 percent 

c. 48.87 percent 

d. 56.01 percent


1
Expert's answer
2021-06-16T09:23:09-0400

0 percent on the first $10,000, 

10 percent on the next $10,000, 

15 percent on the next $10,000 , 

25 percent bin the next $10,000 , and

50 percent on all additional income.

Here Marcus has earns the total income of "\$70,000 (\$60,000 + \$10,000)" . So tax will be calculated on the total amount of tax.

Taxable salary = $70,000 - (5% of Provisional Income tax) - (15.3% of Labor Income) "= \$70,000 - \$3500 - \$9,180 = \$57,320"

"Tax \space on \space \$57,320 = \$10,000 \times 0 \% + \$10,000 \times 10\% + \$10,000 \times 15\% +\$10,000 \times 25\% +\$17,320 \times 50\% = 0 +\$1,000 +\$1,500 + \$2,500 +\$8,660 = \$13,660"

Average Rate "= \frac{\$13,660 }{ \$70,000 }= 17.19\%"

So Option no A is correct answer.